An archaeologist graduate student found a leg bone of a large animal during the building of a new science building . The bone had a carbon-14 decay rate of 14.8 disintegration per minute per gram of carbon .Living organisms have a decay rate of 15.3 disintegrations per minute .How old is the bone? (a) 53.3 years. (b), 25 years. (c), 111years. (d), 83 years. (e), 275 years.
Answer #1

Radioactive decay follows first-order kinetics. By measuring the C-14 activity of a dead sample, its age can be calculated using the expression for radioactive decay as:

t = 2.303/) * log(No/N)

where t is its age, \lambda is the decay constant of C-14, N_{0} is the C-14 activity in a fresh sample of the species, N_{t} is its C-14 activity at present.

Given, N_{t} = 14.8 dpm/g (dpm/g is disintegration per minute per gram)

N_{0} = 15.3 dpm/g

Half-life of carbon-14, t_{1/2} = 5760 years

A = 0.693t12 = 0.693 / (5760 years) = 1.203*10-4 years-1

Substituting these values in the equation t = 2.303/) * log(No/N) ,

Age  t = [ 2.303/(1.203*10-4 years-1) ] * log (15.3 dpm/g / 14.8 dpm/g)

= [ 2.303/(1.203*10-4) ] * log (15.3/14.8) years

= 19143.807 * 0.0144 years

= 275.67 years

Hence the bone is 275 years old.

The correct option is option e - 275 years.